Primary tabs

Story Archive

Category Title Date Excerpt Attached Images
Beam delivery update 24 April 2015 The cyclotron is operational and delivering 40 µA to BL2A and ~65 µA to BL2C4.
Info on new funding for TRIUMF 24 April 2015 News...
Beam delivery update: Cyclotron operational 23 April 2015 Beam was injected shortly after 14:00 this afternoon and we’re currently sitting with 5 µA in BL2A. ISAC target scans have been completed and the plan is to step the current back up to 40 µA as the cyclotron tune and target vacuum allow.
Experimental Result Baartman scores touchdown in magnet design 23 April 2015 Rick Baartman has changed the shapes of magnets used to focus particle beams.
Beam delivery update: Cyclotron RF, electrostatics conditioning 22 April 2015 This morning’s leak-checking went well and the cyclotron has been pumping down all day. RF and inflector/deflector conditioning are already underway and will continue overnight.
Beam delivery update: Problem solved; cyclotron pumping down 21 April 2015 It was determined, on further investigation, that the wire observed to be in the beam plane of the cyclotron was in a location that might be accessible via a roughing port on the cyclotron tank. The Vacuum Group was able to snag the wire using a long tool and pull it underneath a cryopanel and has since started pumping down the cyclotron.  RF conditioning, etc. is expected to take place Thursday with injection to follow on Friday.
Funding Announcements Budget 2015 Renews Support for TRIUMF 21 April 2015 The Government of Canada's Economic Action Plan 2015 commits to providing an additional $45 million in operating funds to TRIUMF over the next five years. TRIUMF applauds the Government’s continued investment in science, technology and innovation.
Beam delivery update: Cyclotron lid to be raised 21 April 2015 No proton beam is expected this week. Exclusion areas will be accessible as a result; any work that might have been planned for upcoming maintenance days should be carried out this week if possible.
Beam delivery update: Maintenance tomorrow 20 April 2015 Tomorrow’s day shift was scheduled as “maintenance if required”; beam will go off at ~06:00 to allow for a couple of hours of cooldown before repairs start around 08:30.
Beam delivery update 17 April 2015 The cyclotron is operational. Beam was injected around 16:30 this afternoon and operators from the Main Control Room are in the process of tuning it through the machine.

Pages

Subscribe to Recent Stories